OpenStudy (anonymous):

evalutate log_7 root 343

jimthompson5910 (jim_thompson5910):

Hint: sqrt(343) = sqrt(7^3) sqrt(343) = (7^3)^(1/2) sqrt(343) = 7^(3/2)